Explain the meaning of the term “variable”. Write a puzzle for a 10-year child. Write
an equation with a variable in it which gives a solution to the question raised by the
puzzle. Also give a solution to the puzzle and interpret it
1
Expert's answer
2020-04-03T14:59:44-0400

In mathematics, a variable is a symbol used to represent an arbitrary element of a set.

[https://en.m.wikipedia.org/wiki/Variable_(mathematics)]

Let's imagine that we walk every day during one week (7 days). We know that every 1 minute of walking involves 100 steps. So, we need to determine the number of steps. We also know that on the first day we walked 30 minutes, on the 2nd day - 10 steps, on the 3rd day - 15 steps, on the 4th day - 20 steps, on the 5th day - 10 steps, on the 6th day - 5 steps, on the 7th day - 30 steps.

The function we will use is f(x) =y*100, where f is a variable (our steps), y is time.

So, know we can solve the equation and find our steps every day (our variable).

1st f(x) =30*100=3000 steps

2nd f(x) =10*100=1000 steps

3rd f(x) =15*100=1500 steps

4th f(x) =20*100=2000 steps

5th f(x) =10*100=1000 steps

6th f(x) =5*100=500 steps

7th f(x) =30*100=3000 steps

So, we can say that the variable is something that can change depending on the measure we take.